I am not familiar with this kind of questions. Please help me solve this. Your work will be greatly appreciated. THX !!
Answer #1

Since X is a continuous random variable and P(X<0) = 0. We have

f_X(x) = \frac{\mathrm{d}(F_X(x)) }{\mathrm{d} x}\Rightarrow P(X\leq x) = F(x) = \int_{0}^{x}f(x)dx

a) Since 0 < b < a. Now

P(a \leq X\leq b) = \int_{a}^{b}f(x)dx = \int_{a}^{b}\frac{\mathrm{d} F(x)}{\mathrm{d} x}dx=\\ \int_{a}^{b}\mathrm{d} F(x) = F(b)-F(a)

F_X(0) = P(X\leq 0) = 0\text{ (given in question)}

b) a > 0.

\int_{0}^{a}xf_X(x)dx = x\int_{0}^{a}f_X(x)dx-\int_{0}^{a}[\frac{\mathrm{d} (x)}{\mathrm{d} x}\int f_X(x)dx]dx=\\ xF_X(x)|_{0}^{a} - \int_{0}^{a} F_X(x)dx = (aF_X(a)-\lim_{x\rightarrow 0}xF_X(x)) - \int_{0}^{a} F_X(x)dx.\\ Now,\: \: \lim_{x\rightarrow 0}xF_X(x) = \text{(By L'Hospital Rule)} \lim_{x\rightarrow 0}f_X(x) = 0.\\ Therefore,\: \: \int_{0}^{a}xf_X(x)dx =aF_X(a) - \int_{0}^{a} F_X(x)dx= \int_{0}^{a} F_X(a)dx-\int_{0}^{a} F_X(x)dx =\\ \int_{0}^{a} (F_X(a) - F_X(x))dx

c)

E[X] = \int_{0}^{\infty}xf_X(x)dx= \int_{0}^{\infty} (F_X(\infty) - F_X(x))dx.\\ Now, \: \: \lim_{x \rightarrow \infty}F_X(x) = F_X(\infty) = P(X\leq \infty) = 1(\text{ The number is always less than infinity} ).\\ Therefore, \: \: E[X] = \int_{0}^{\infty} (1 - F_X(x))dx.\\ Also, \: \: P(X > x) = 1-P(X\leq x) = 1 - F_X(x). \\ Hence, \: \: E[X] = \int_{0}^{\infty} P(X>x)dx